OpenStudy (anonymous):

Solve for x: |2x + 6| - 4 = 20

OpenStudy (anonymous):

x : |2x + 6| - 4 = 20 +4 +4 x: |2x+6| = 24 -6 -6 x: 2x = 18 x/2 = 18/2 Answer x= 9

OpenStudy (amistre64):

there will generally be 2 answers for s with absolute values: in this case we want to know when: 2x+6 = 24 and 2x+6 = -24